OpenStudy (anonymous):

Solve x^3+9x^2+8x-60=0 Please show work thx

OpenStudy (tanya123):

Simplifying x3 + -9x2 + 8x + 60 = 0 Reorder the terms: 60 + 8x + -9x2 + x3 = 0 Solving 60 + 8x + -9x2 + x3 = 0 Solving for variable 'x'. The solution to this equation could not be determined.

OpenStudy (raden):

x^3+9x^2+8x-60=0 start from find the factors of 60, are {+-1, +-2, +-3, +-4, +-5, +-12, +-15, +-20, +- 30, +-60} let one of the solution is 2, check it : x^3+9x^2+8x-60=0 (2)^3 + 9(2)^2 + 8(2) - 60 = 0 8 + 36 + 16 - 60 = 0 0 = 0 (TRUE) then you can use long division or horner's diagram to get the other solutions

OpenStudy (mathmate):

or by grouping: x^3+9x^2+8x-60 =(x^3-2x^2) +(11x^2-22x) +(30x-60) =(x-2)(x^2+11x+30) =(x-2)(x+6)(x+5)